# Unsupervised Machine Learning for Classifying CHIME Fast Radio Bursts and Investigating Empirical Relations

### v1

@@ 3 Method
Dimensionality reduction and clustering are two types of unsupervised machine learning methods used in this paper. First, we apply the dimensionality reduction algorithm to automatically convert high-dimensional data into low-dimensional data. Then, we use a clustering algorithm to classify the reduced-dimensional data based on their similarities.
@@ 3.1.1 Dimensionality Reduction
The use of UMAP has been extensively explored in many studies. For FRB classification, the three parameters n_compnent, n_neighbors, and min_dist have a more significant impact on the classification results. Meanwhile, we also experimented with modifying other parameters and found that they had minimal effect on the classification outcomes. Therefore, in this study, we choose to focus on adjusting these three parameters of UMAP. n_compnent allows us to determine the dimensionality of the reduced space where the data will be embedded. In our work, we set $\mathtt{n\_compnent}=2$ for all features, projecting the data onto a 2D plane for visual representation.n_neighbors controls how UMAP balance the local and globol structure of data. UMAP achieves this by controlling the size of the local neighborhood it considers when attempting to learn the underlying structure of the data. This means that low values of n_neighbors will cause UMAP to focus on very local structures, potentially at the cost of missing the overall global structure. On the other hand, higher values of n_neighbors will push UMAP to consider larger neighborhoods around each point, capturing the broader structure of the data, but possibly losing finer details. min_dist decides how tightly UMAP can pack points together in low-dimensional space. Basically, it sets the minimum distance between points in the low-dimensional space. Lower values of min_dist will lead to more tightly packed, “clumpier” embeddings, which can be beneficial for identifying clusters or preserving finer topological details. In contrast, higher values of min_dist will prevent the points from being tightly packed, instead focusing on maintaining the broader topological structure. We scan n_neighbors from 2 to 50 and min_dist from 0.0 to 0.99. In this paper, we take n_neighbors = 21, and min_dist = 0.03.
@@ 3.1.2 Clustering Algorithms
In this work, we used k-means (MacQueen et al., 1967; Lloyd, 1982) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN Campello et al., 2013, 2015; McInnes et al., 2017) to group the reduced-dimensional data into different clusters. K-means clustering is based on the distance between each data point and its corresponding cluster center. It aims to minimize the distance between the points and their respective centers, effectively grouping similar points into clusters based on proximity in the feature space. Initially, the k-means algorithm selects k random points as the initial cluster centers, calculates the Euclidean distance of each data point from these centers, and assigns each point to the nearest center. Then, it recalculates the mean of each cluster, updating the cluster centers. This process is repeated iteratively until the cluster centers stabilize, minimizing the overall variance within the cluster (for detals, see Fotopoulou, 2024, and reference their in). We use sklearn.cluster.KMeans [^4] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html to perform the k-means clustering algorethm. The essential hyperparameter is n_clusters, it means how many clusters are present in the model.
@@
Figure 3: The results of dimensionality reduction and clustering. Left panel: The distribution of UMAP-embedded data. Blue dots represent non-repeaters, while green and red crosses indicate repeaters from Cat1 and Cat2023, respectively. Middle panel: The clustring result of UMAP-embedded data from k-measn. Cluster 0, 1, and 2 are represented by blue dots, orange stars, and red crosses, respectively. Right panel: The clustring result of UMAP-embedded data from HDBSCAN. Gray triangles, blue dots, royal blue squares, orange stars, tomato diamonds, and red crosses represent Noise, and Cluster 0, 1, 2, 3, and 4, respectively.
Algorithm Cluster Non-repeater number Repeater number Repeater candidate number Total number Repeater precentage Recall Repeating rate
K-means Cluster 0 210 0 0 210 0% 100% 61.7%
Cluster 1 0 98 196 294 33.3%
Cluster 2 0 132 103 235 56.2%
HDBSCAN Noise 45 14 0 59 23.7% 93.5% 37.0%
Cluster 0 169 0 0 169 0%
Cluster 1 138 15 0 153 9.8%
Cluster 2 0 88 118 206 42.7%
Cluster 3 0 66 28 94 70.2%
Cluster 4 0 47 11 58 81.0%
To improve the clustering results, we optimize the hyperparameters of the clustering algorithm by maximizing the mean silhouette coefficient  [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.silhouette_score.html (Rousseeuw, 1987) across all samples. The silhouette coefficient measures how well a sample fits within its assigned cluster, with values ranging from -1 to 1. Higher values indicate more tightly grouped, well-defined clusters. We show the mean silhouette coefficients changed with the hyperparameters of the clustering algorithm in Figure 2. In this work, we chose the hyperparameters with max mean silhouette coefficients, which is n_clusters = 3 for k-means, min_cluster_size = 37  [^7] In fact, when we fixed min_samples at 3, the mean silhouette coefficients remained constant as min_cluster_size varied from 37 to 60. However, the number of noise points increased during this range. Therefore, we chose min_cluster_size = 37 for this study. and min_samples = 3 for HDBSCAN.
@@ 4.1 Dimensionality reduction and clustering
Then we used two types of clustering algorithm, k-means and HDBSCAN, to cluster the two-dimensional UMAP embedding. We present the clustering results in the middle and right panels of Figure 3, based on the hyperparameters of k-means and HDBSCAN discussed in Section 3.1.2. If the proportion of repeaters in a cluster exceeds 30%, we classify it as a “ repeater cluster ”, with the non-repeaters within it considered as “ repeater candidates ”. Conversely, if the proportion is below 30%, it is classified as an “ non-repeater cluster ”.
@@
Figure 5: Same as Figure 4, but displays the results from HDBSCAN. The blue, royal blue, orange, tomato, and red histogram step lines represent Cluster 0, 1, 2, 3, and 4, respectively.
@@
As shown in the middle panel of Figure 3, the k-means algorithm divided the UMAP-embedded data into three clusters. Cluster 0 contains 210 FRBs, all of which are non-repeaters. Cluster 2 consists of 294 FRBs, including 98 repeaters and 196 repeater candidates (the percentage of repeaters is 33.3%). Cluster 3 contains 235 FRBs, with 132 repeaters (the percentage of repeaters is 56.2%) and 103 repeater candidates. All repeaters are classified into repeater clusters, giving a recall of 100%. Out of the 509 non-repeaters, 299 repeater candidates were identified (from 269 non-repeater sources). If these repeater candidates are real, the repeating rate of FRBs would reach approximately 61.7%, exceeding the predicted rates from studies such as Chen et al. (2022), Zhu-Ge et al. (2022), and Yang et al. (2023).
Figure 6: The distribution of non-repeaters (blue dots), repeater candidates (orange diamonds), and repeaters (red crosses) is presented in the clustered UMAP-embedded data from both k-means (left panel) and HDBSCAN (right panel). And the gray triangles are represent noises. The green stars mark the first bursts of six repeaters that were misidentified as non-repeaters. The black circles indicate the localized non-repeaters.
@@
As shown in the right panel of Figure 3, HDBSCAN divided the UMAP-embedded data into five clusters and a noise cluster. Cluster 0 contains 169 FRBs, all of which are non-repeaters. Cluster 1 includes 15 repeaters and 138 non-repeaters, with repeaters making up only 9.8%. Clusters 2, 3, and 4 contain 206, 94, and 58 FRBs, with 88, 66, and 47 repeaters (repeaters accounting for 42. 7%, 70. 2%, and 81. 0%, respectively) and repeater candidates numbering 118, 28, and 11 (a total of 157 candidates, corresponding to 141 non-repeater sources). The recall for the 739 samples is 93.1%. The overall repeating rate is around 37.9%, slightly lower than the results of Yamasaki et al. (2023); McGregor & Lorimer (2024), but comparable to those of Chen et al. (2022), Zhu-Ge et al. (2022), and Yang et al. (2023). Detailed clustering results for both algorithms are also presented in Table 1.
@@
We plot the feature distributions of different clusters generated by k-means and HDBSCAN in Figure 4 and Figure 5, respectively. As shown in Figure 4, the rest-frame frequency width differs the most across clusters, indicating that repeater clusters tend to have narrower frequency bandwidths. Significant differences are also observed in the distributions of the spectral index, highest frequency, redshift, energy, luminosity, and brightness temperature among the clusters. These results are consistent with the feature distribution of repeaters and non-repeaters shown in Figure 1. In Figure 5, almost all features show notable differences in their distributions between clusters, with rest-frame frequency width once again being the most distinct, similar to the results in Figure 4.
@@
Figure 7: The empirical relationships between scattering time (${\rm log\,}\Delta t_{sc}$) and rest-frame width (${\rm log\,}\Delta t_{rw}$) (left panel), and between spectral acceleration ($r$) and spectral index ($\gamma$) (right panel) are shown for FRBs in different clusters from k-means, represented by blue, orange, and red lines, respectively. The corresponding $R^{2}$ values (in blue, orange, and red) are also provided. See Section 4.2 for details.
        x         y         cluster         $a$         $b$         $R^{2}$
@@
4.2 Empirical Relationships
In this paper, we further analyze the potential two-dimensional empirical relations within different clusters identified by various clustering algorithms. We pair the 16 parameters of FRBs from different clusters and linearly fit the data points using scipy.stats.linregress [^8] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.linregress.html, which performs linear least-squares regression. The form of the two-dimensional empirical relationship is given as $y=a\,x+b$, and the goodness of fit is evaluated using the score (coefficient of determination), defined as $R^{2}=1-\sum_{i}(y_{i}-\hat{y}_{i})^{2}/\sum{i}(y_{i}-\bar{y}_{i})^{2}$, where $y_{i}$, $\hat{y}_{i}$, and $\bar{y}_{i}$ are the observed values, the regressed values and the mean of the observed values, respectively. The closer $R^{2}$ is to 1, the better the model fits the data.
Figure 8: The empirical relationships between scattering time (${\rm log\,}\Delta t_{sc}$) and brightness temperature (${\rm log\,}T_{B}$) (left panel), and between spectral acceleration ($r$) and spectral index ($\gamma$) (right panel) are shown for FRBs in different clusters from HDBSCAN, represented by blue, royal blue, orange, tomato, and red lines, respectively. The corresponding $R^{2}$ values (in blue, royal blue, orange, tomato, and red) are also provided. See Section 4.2 for details.
         x          y          cluster          $a$          $b$          $R^{2}$
@@ 4.2 Empirical Relationships
We set a high score threshold of $R^{2}>0.5$ to filter out well-fitted empirical relations and excluded parameter combinations that inherently have a linear relationship (e.g., luminosity and flux). For the clusters from k-means, the selected empirical relations are shown in Figure 7, and the slope, intercept, and score of the empirical relations are listed in Table 2. The left panel shows the empirical relation between two independent parameters: scattering time (${\rm log\,}\Delta t_{sc}$) and rest-frame width (${\rm log\,}\Delta t_{rw}$). We observe that $R^{2}>0.5$ only in cluster 2, corresponding to a repeater cluster, whereas the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation appears less significant in the other two clusters. However, the slope and intercept of the relation ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ are similar in all three clusters. The right panel displays the empirical relation between two other independent parameters: spectral acceleration ($r$) and spectral index ($\gamma$). In contrast to the ${\rm log\,}\Delta t_{rw}-{\rm log\,}\Delta t_{sc}$ relation, the $r-\gamma$ relation is less evident in cluster 2, but is very pronounced in clusters 0 and 1 ($R^{2}>0.8$), with similar slopes and intercepts.
@@
Figure 8 shows the empirical relations for different clusters from HDBSCAN, and Table 3 lists the slope, intercept, and score of these empirical relations. The left panel displays the relation between rest-frame width (${\rm log\,}\Delta t_{sc}$) and the brightness temperature (${\rm log\,}T_{B}$), which are also two independent parameters. We can see that only cluster 0 (a non-repeater cluster) has a significant ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation with $R^{2}>0.6$, while the other clusters show no strong ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation. The right panel shows the empirical relation between spectral acceleration ($r$) and spectral index ($\gamma$). All clusters exhibit a significant $r-\gamma$ relation, especially clusters 0-3, where $R^{2}>0.8$. Notably, clusters 0 and 3 (a repeater cluster) have very similar slopes and intercepts for the $r-\gamma$ relation, suggesting that certain repeaters might have spectral properties similar to those of non-repeaters, a trend also observed in the k-means results..

### v2

@@ 3 Method
Dimensionality reduction and clustering are two types of unsupervised machine learning methods used in this paper. First, we apply the dimensionality reduction algorithm to automatically convert high-dimensional data into low-dimensional data. Then, we use a clustering algorithm to cluster the reduced-dimensional data based on their similarities.
@@ 3.1.1 Dimensionality Reduction
The use of UMAP has been extensively explored in many studies (e.g. Zhu-Ge et al., 2022; Yang et al., 2023; Raquel et al., 2023; Chen et al., 2023). For FRB classification, the three parameters n_components, n_neighbors, and min_dist have a more significant impact on the classification results. Meanwhile, we also experimented with modifying other parameters and found that they had minimal effect on the classification outcomes. Therefore, in this study, we choose to focus on adjusting these three parameters of UMAP. n_components allows us to determine the dimensionality of the reduced space where the data will be embedded. In our work, we set $\mathtt{n\_components}=2$ for all features, projecting the data onto a 2D plane for visual representation. n_neighbors controls how UMAP balances the local and global structure of data. UMAP achieves this by controlling the size of the local neighborhood it considers when attempting to learn the underlying structure of the data. This means that low values of n_neighbors will cause UMAP to focus on very local structures, potentially at the cost of missing the overall global structure. On the other hand, higher values of n_neighbors will push UMAP to consider larger neighborhoods around each point, capturing the broader structure of the data, but possibly losing finer details. min_dist decides how tightly UMAP can pack points together in low-dimensional space. Basically, it sets the minimum distance between points in the low-dimensional space. Lower values of min_dist will lead to more tightly packed, “clumpier” embeddings, which can be beneficial for identifying clusters or preserving finer topological details. In contrast, higher values of min_dist will prevent the points from being tightly packed, instead focusing on maintaining the broader topological structure. We scan n_neighbors from 2 to 50 and min_dist from 0.0 to 0.99. In this paper, we take n_neighbors = 21, and min_dist = 0.03.
@@ 3.1.2 Clustering Algorithms
In this work, we used k-means (MacQueen et al., 1967; Lloyd, 1982) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN, Campello et al., 2013, 2015; McInnes et al., 2017) to group the reduced-dimensional data into different clusters. K-means clustering is based on the distance between each data point and its corresponding cluster center. It aims to minimize the distance between the points and their respective centers, effectively grouping similar points into clusters based on proximity in the feature space. Initially, the k-means algorithm selects k random points as the initial cluster centers, calculates the Euclidean distance of each data point from these centers, and assigns each point to the nearest center. Then, it recalculates the mean of each cluster, updating the cluster centers. This process is repeated iteratively until the cluster centers stabilize, minimizing the overall variance within the cluster (for details, see Fotopoulou, 2024, and references therein). We use sklearn.cluster.KMeans [^4] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html to perform the k-means clustering algorithm. The essential hyperparameter is n_clusters which means how many clusters are present in the model.
@@
To improve the clustering results, we optimize the hyperparameters of the clustering algorithm by maximizing the mean silhouette coefficient  [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.silhouette_score.html (Rousseeuw, 1987) across all samples. The silhouette coefficient measures how well a sample fits within its assigned cluster, with values ranging from -1 to 1. Higher values indicate more tightly grouped, well-defined clusters. We show the mean silhouette coefficients changed with the hyperparameters of the clustering algorithm in Figure 2. In this work, we chose the hyperparameters with max mean silhouette coefficients, which are n_clusters = 3 for k-means, min_cluster_size = 37  [^7] In fact, when we fixed min_samples at 3, the mean silhouette coefficients remained constant as min_cluster_size varied from 37 to 60. However, the number of noise points increased during this range. Therefore, we chose min_cluster_size = 37 for this study. and min_samples = 3 for HDBSCAN. We also include the condensed tree of HDBSCAN in Appendix A to highlight the persistence and stability of the clustering results.
@@ 4.1 Dimensionality reduction and clustering
Figure 3: The results of dimensionality reduction and clustering. Left panel: The distribution of UMAP-embedded data. Blue dots represent non-repeaters, while green and red crosses indicate repeaters from Cat1 and Cat2023, respectively. Middle panel: The clustering result of UMAP-embedded data from k-means. Cluster 0, 1, and 2 are represented by blue dots, orange stars, and red crosses, respectively. Right panel: The clustering result of UMAP-embedded data from HDBSCAN. Gray triangles, blue dots, cyan squares, orange stars, light coral diamonds, and red crosses represent Noise, and Cluster 0, 1, 2, 3, and 4, respectively.
Algorithm Clusters Non-repeater number Repeater number Repeater candidate number Total number Repeater burst percentage Recall Repeater source percentage
K-means 0 210 0 0 210 0% 100% 61.7%
1 0 98 196 294 33.3%
2 0 132 103 235 56.2%
HDBSCAN Noise 45 14 0 59 23.7% 93.5% 37.0%
0 169 0 0 169 0%
1 138 15 0 153 9.8%
2 0 88 118 206 42.7%
3 0 66 28 94 70.2%
4 0 47 11 58 81.0%
@@
Figure 5: Same as Figure 4, but displays the results from HDBSCAN. The blue, cyan, orange, light coral, and red histogram step lines represent Cluster 0, 1, 2, 3, and 4, respectively.
Then we used two types of clustering algorithms, k-means and HDBSCAN, to cluster the two-dimensional UMAP embedding. We present the clustering results in the middle and right panels of Figure 3, based on the hyperparameters of k-means and HDBSCAN discussed in Section 3.1.2. If the proportion of repeaters in a cluster exceeds 30%, we classify it as a “ repeater cluster ”, with the non-repeaters within it considered as “ repeater candidates ”. Conversely, if the proportion is below 30%, it is classified as an “ non-repeater cluster ”.
@@
As shown in the middle panel of Figure 3, the k-means algorithm divided the UMAP-embedded data into three clusters. Cluster 0 contains 210 FRBs, all of which are non-repeaters. Cluster 2 consists of 294 FRBs, including 98 repeaters and 196 repeater candidates (the repeater burst percentage is 33.3%). Cluster 3 contains 235 FRBs, with 132 repeaters (the repeater burst percentage is 56.2%) and 103 repeater candidates. All repeaters are classified into repeater clusters, giving a recall of 100%. Out of the 509 non-repeaters, 299 repeater candidates were identified (from 269 non-repeater sources). If these repeater candidates are real, the repeater source percentage of FRBs would reach approximately 61.7%, exceeding the predicted rates from studies such as Chen et al. (2022), Zhu-Ge et al. (2022), and Yang et al. (2023).
@@
As shown in the right panel of Figure 3, HDBSCAN divided the UMAP-embedded data into five clusters and a noise cluster. Cluster 0 contains 169 FRBs, all of which are non-repeaters. Cluster 1 includes 15 repeaters and 138 non-repeaters, with repeater bursts making up only 9.8%. Clusters 2, 3, and 4 contain 206, 94, and 58 FRBs, with 88, 66, and 47 repeaters (repeater bursts percentage accounting for 42.7%, 70.2%, and 81.0%, respectively) and repeater candidates numbering 118, 28, and 11 (a total of 157 candidates, corresponding to 141 non-repeater sources). The recall for the 739 samples is 93.1%. The overall repeater source percentage is around 37.9%, slightly lower than the results of Yamasaki et al. (2023); McGregor & Lorimer (2024), but comparable to those of Chen et al. (2022), Zhu-Ge et al. (2022), and Yang et al. (2023). Detailed clustering results for both algorithms are also presented in Table 1.
@@
We plot the feature distributions of different clusters generated by k-means and HDBSCAN in Figure 4 and Figure 5, respectively. As shown in Figure 4, the rest-frame frequency width differs the most across clusters, indicating that repeater clusters tend to have narrower frequency bandwidths. Significant differences are also observed in the distributions of the spectral index, highest frequency, redshift, energy, luminosity, and brightness temperature among the clusters. These results are consistent with the feature distribution of repeaters and non-repeaters shown in Figure 1. In Figure 5, almost all features show notable differences in their distributions among clusters, with rest-frame frequency width once again being the most distinct, similar to the results in Figure 4.
Figure 6: The distribution of non-repeaters (blue dots), repeater candidates (orange diamonds), and repeaters (red crosses) is presented in the clustered UMAP-embedded data from both k-means (left panel) and HDBSCAN (right panel). The gray triangles represent noises. The green stars mark the first bursts of six repeaters that were misidentified as non-repeaters in Cat1 but are later identified as repeaters in Cat2023. The black circles indicate the localized non-repeaters.
@@
Figure 7: The empirical relationships between scattering time (${\rm log\,}\Delta t_{sc}$) and rest-frame width (${\rm log\,}\Delta t_{rw}$) (left panel), and between spectral acceleration ($r$) and spectral index ($\gamma$) (right panel) are shown for FRBs in different clusters from k-means. These relationships are represented by blue, orange, and red lines, corresponding to clusters 0, 1, and 2, respectively. The associated $R^{2}$ values are shown in the same colors. Additionally, the dashed magenta line and its corresponding $R^{2}$ represent the combined data from clusters 1 and 2. See Section 4.2 for details.
        x         y         Clusters         $a$         $b$         $R^{2}$
@@
Clusters Relation $p$-value ${\rm log\,}\Delta t_{sc}\,-\,{\rm log\,}\Delta t_{rw}$ $r\,-\,\gamma$
0 & 1 0.0016 0.5177
0 & 2 1.75e-06 6.28e-05
1 & 2 0.1975 1.11e-16
0 & [1, 2] 2.49e-06 0.0190
Table 3: The $p$-values of Chow tests for k-means clusters.
Regarding FRB 20190226B, which was misidentified only by HDBSCAN, we analyzed its 16 parameters and found no distinct characteristics suggesting it is a non-repeating burst.
@@
Figure 8: The empirical relationships between scattering time (${\rm log\,}\Delta t_{sc}$) and brightness temperature (${\rm log\,}T_{B}$) (left panel), and between spectral acceleration ($r$) and spectral index ($\gamma$) (right panel) are shown for FRBs in different clusters from HDBSCAN. These relationships are represented by blue, cyan, orange, light coral, and red lines, corresponding to clusters 0, 1, 2, 3, and 4, respectively. The associated $R^{2}$ values are shown in the same colors. Additionally, the dashed black (magenta) line and its corresponding $R^{2}$ represent the combined data from clusters 0 and 1 (2, 3, and 4). See Section 4.2 for details.
         x          y          Clusters          $a$          $b$          $R^{2}$
@@
Clusters Relation $p$-value ${\rm log\,}\Delta t_{sc}\,-\,{\rm log\,}T_{B}$ $r\,-\,\gamma$
0 & 1 0.0041 1.13e-10
0 & 2 0.1701 1.75e-09
0 & 3 0.0483 0.4561
0 & 4 0.5232 1.11e-16
1 & 2 0.2403 0.0019
1 & 3 8.12e-07 6.34e-12
1 & 4 0.0014 1.11e-16
2 & 3 0.0002 1.11e-16
2 & 4 0.0379 1.11e-16
3 & 4 0.6024 1.11e-16
[0, 1] & [2, 3, 4] 2.43e-11 1.97e-08
Table 5: The $p$-values of Chow tests for HDBSCAN clusters.
4.2 Empirical Relationships
In this paper, we further analyze the potential two-dimensional empirical relations within different clusters identified by various clustering algorithms. We pair the 16 parameters of FRBs from different clusters and linearly fit the data points using scipy.stats.linregress [^8] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.linregress.html, which performs linear least-squares regression. The form of the two-dimensional empirical relationship is given as $y=a\,x+b$, and the goodness of fit is evaluated using the score (coefficient of determination), defined as $R^{2}=1-\sum_{i}(y_{i}-\hat{y}_{i})^{2}/\sum_{i}(y_{i}-\bar{y}_{i})^{2}$, where $y_{i}$, $\hat{y}_{i}$, and $\bar{y}_{i}$ are the observed values, the regressed values and the mean of the observed values, respectively. The closer $R^{2}$ is to 1, the better the model fits the data. To quantify the statistical differences in the slopes and intercepts of empirical relationships across different clusters, we implemented the Chow test for evaluation (Chow, 1960). The null hypothesis of the Chow test is that the two samples come from the same regression model. If the $p$-value is less than 0.05, the null hypothesis is rejected, indicating that the two samples come from different regression models.
@@ 4.2 Empirical Relationships
We set a high score threshold of $R^{2}>0.5$ to filter out well-fitted empirical relations and excluded parameter combinations that inherently have a linear relationship (e.g., luminosity and flux). For the clusters from k-means, the selected empirical relations are shown in Figure 7. The slope, intercept, and score of the empirical relations are listed in Table 2. The $p$-values of the Chow test between different clusters are listed in Table 3. The left panel of Figure 7 shows the empirical relation between two independent parameters: scattering time (${\rm log\,}\Delta t_{sc}$) and rest-frame width (${\rm log\,}\Delta t_{rw}$). We observe that $R^{2}>0.5$ only in cluster 2, corresponding to a repeater cluster, whereas the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation appears less significant in the other two clusters. However, the slope and intercept of the relation ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ are similar in clusters 1 and 2. Similarly, the $p$-value of the Chow test between clusters 1 and 2 is greater than 0.05, indicating that they likely come from the same regression model. This may be because both clusters are repeater clusters. The right panel of Figure 7 displays the empirical relation between two other independent parameters: spectral acceleration ($r$) and spectral index ($\gamma$). In contrast to the ${\rm log\,}\Delta t_{rw}-{\rm log\,}\Delta t_{sc}$ relation, the $r-\gamma$ relation is less evident in cluster 2, but is very pronounced in clusters 0 and 1 ($R^{2}>0.8$), with similar slopes and intercepts. Moreover, the $p$-value of the Chow test between these two clusters even reaches around 0.5, indicating that some repeaters may share spectral properties with non-repeaters. Combining the two repeater clusters, cluster 1 and cluster 2, the relation ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ yields an $R^{2}$ value of 0.4608. Although this is less than 0.5, it still exhibits a distinct empirical correlation compared to non-repeating bursts. The $r-\gamma$ relation for the combined clusters has an $R^{2}$ value of 0.3767, which is even lower than that of cluster 2. This may result from the substantial disparity between the $r-\gamma$ relations of cluster 1 and cluster 2, as well as the relatively weak correlation observed in cluster 2. The $p$-values of the Chow test between the non-repeater cluster (Cluster 0) and the repeater clusters (Clusters 1 and 2) are all below 0.05 for both empirical relations. This result strengthens the idea that FRBs can be classified into repeaters and non-repeaters.
@@
Figure 8 shows the empirical relations for different clusters from HDBSCAN, and Table 4 lists the slope, intercept, and score of these empirical relations. The left panel displays the relation between rest-frame width (${\rm log\,}\Delta t_{sc}$) and the brightness temperature (${\rm log\,}T_{B}$), which are also two independent parameters. We can see that only cluster 0 (a non-repeater cluster) has a significant ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation with $R^{2}>0.6$, while the other clusters show no strong ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation. Even if the $p$-values of the Chow test between some clusters are greater than 0.05, it does not necessarily indicate that they originate from the same regression model. The right panel shows the empirical relation between spectral acceleration ($r$) and spectral index ($\gamma$). All clusters exhibit a significant $r-\gamma$ relation, especially clusters 0-3, where $R^{2}>0.8$. Notably, clusters 0 and 3 (a repeater cluster) have very similar slopes and intercepts for the $r-\gamma$ relation (the $p$-value of the Chow test is 0.45), suggesting that certain repeaters might have spectral properties similar to those of non-repeaters, a trend also observed in the k-means results. We also analyzed the relations using the combined data from clusters 0 and 1 (non-repeater clusters), as well as from clusters 2, 3, and 4 (repeater clusters). For the ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation, the $R^{2}$ value for the non-repeater clusters is less than 0.5, likely due to the scattered data in cluster 1. However, it still performs better than the repeater cluster, which has an $R^{2}$ value of 0.0444, indicating that the ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation is virtually non-existent for the repeater cluster. The $r-\gamma$ relation is still apparent in the non-repeater clusters, but the $R^{2}$ value is less than 0.5 in the repeater clusters. This may be due to the significant differences in slope and intercept of the $r-\gamma$ relation across clusters 2, 3, and 4. The $p$-values of the Chow test between non-repeater clusters and repeater clusters are all significantly below 0.05 for both empirical relations, aligning with the results from k-means clustering.